Y=2/3x and Y=-2/7x+4/7
I am not even sure how to set this up. Do I need to have the y-intercept for the first system of equation befor I can solve? Can you please help me set this problem up? Thanks for your help!

Answer by phillv(13)

You do not need the Y intercept.
Just substitute 2/3x for Y in the second equation.
This gives us 2/3x=-2/7x+4/7
2/3x=(4-2x)/7
multiply both sides by 7
14/3x=4-2x
multiply both sides by 3
14x=12-6x
simplify
20x=12
x=12/20=3/5
x=3/5
substitute 3/5 for x in either of the original equations
Y=(2/3)(3/5)
Y=2/5
X=3/5 and Y=2/5 will satisfy both equations.